### v1

# Displacement Memory Effect from Supersymmetry

Abstract
We explain the recent results regarding the displacement memory effect (DME) of plane gravitational waves using supersymmetric quantum mechanics. This novel approach stems from the fact that both geodesic and Schrödinger equations are Sturm-Liouville boundary value problems. Supersymmetry provides a unified framework for Pöschl-Teller and Scarf profiles and restores the critical values of wave amplitudes for DME in a natural way. Within our framework, we obtain a compact formula for DME in terms of the asymptotic values of the superpotential and the geodesics. In addition, this new technique enables us to build plane and gravitational waves with 2-transverse directions using superpartner potentials. Lastly, we study DME within a singular wave profile inspired by supersymmetric quantum mechanics, which shows the broader applicability of our method.
Keywords: Displacement memory effect, plane and gravitational waves, supersymmetric quantum mechanics, Sturm-Liouville problem.
1 Introduction
The memory effect of plane waves can be mainly divided into two: the displacement memory effect (DME) [1] and the velocity memory effect (VME) [2, 3, 4]. These effects are of interest as they can be captured by future gravitational wave detectors. In this paper, we address the problem of DME of plane and gravitational waves (GWs) with a novel perspective using supersymmetric quantum mechanics (SUSY QM) [5, 6]. Our main motivation is to investigate the recent results showing the existence of DME in Pöschl-Teller and Scarf GW profiles in the light of SUSY QM.
Previous studies (see [7, 8] and references therein) suggested that plane waves might only have a VME: after the passing of a short burst of a gravitational wave, transverse geodesics have a non-vanishing constant velocity. Only very recently, it has been reported that plane GWs, under special conditions, also can exhibit a DME: permanent displacement with vanishing velocity. Namely, DME occurs only in particular setups with fine-tuned parameters [9, 10, 11] (see also [12]).
Our starting point is the fact that geodesics having DME conditions do satisfy a Sturm-Liouville problem with Neumann boundary conditions. This leads to an equation that we are already familiar with from non-relativistic QM. Therefore, DME of plane GWs and the related conditions can be sought for by using the Schrödinger theory, and the desired geodesics, in principle, could be found analytically by the methods of SUSY QM [5, 6].
This common mathematical origin of geodesics of plane waves and quantum mechanics sheds light on the recent results of those half-waves and magical numbers of the former by relating it to the bound states and quantum numbers of the latter. We show that SUSY elegantly and simply produces these particular setups. We also borrow a singular potential from SUSY QM and discuss the possibility of finding DME within singular GWs, such as the Lukash GW studied in cosmological models.
Plane waves, being models for gravitational radiation, are already significant spacetimes on their own [13]. They are Penrose limits of any spacetime [14], and lower dimensional mechanics can be embedded into them via the Eisenhart lift [15, 16, 17].
As a novelty, we now show that the necessary conditions for their DME relate them to integrable potentials in SUSY QM. Using this interrelation, we obtain the geodesics with DME out of SUSY methods in a compact and beautiful manner. SUSY also helps us to understand the conditions to build exact GWs exhibiting DME in both directions.
Let us note the work of C. Duval et al. [18], who discussed Schrödinger theory in the Bargmann framework. The geodesic equations of spinning particles in gravitational backgrounds were worked out using supersymmetry in [19]. Our study differs from those both in its content and in its scope.
The outline of this paper is the following: In Section 2, we introduce our main motivation, which is DME of the plane waves. Then, we apply SUSY methods of quantum mechanics to plane wave geodesics and discuss how DME occurs in Section 3. 4-dimensional plane waves and their DME are discussed in Section 4. In Section 5, we commented on the DME of singular wave profiles. In our last Section, Section 6, we summarize our results, add our comments, and ask further questions. For completeness, we provide the necessary tools of SUSY QM in the Appendix A.
2 Plane waves and memory effects
In standard Brinkmann coordinates, a generic $d+2$-dimensional plane wave metric is given as
$ds^{2}=\delta_{ij}dX^{i}dX^{j}+2dUdV+K_{ij}(U)X^{i}X^{j}dU^{2},$ (2.1)
where $U,V$ are light-cone coordinates and $X^{i}$ with $i=1,...,d$, are transverse coordinates.
To discuss DME and VME, we consider geodesic motion. Plane waves (2.1) have a covariantly constant null Killing vector $\frac{\partial}{\partial V}$. Therefore, the coordinate $U$ can be chosen as an affine parameter, and geodesic equations for the remaining coordinates turn out to be
$\frac{d^{2}X^{i}}{dU^{2}}-K_{ij}X^{j}=0,$ (2.2a)
$\frac{d^{2}V}{dU^{2}}+\frac{1}{2}\frac{dK_{ij}}{dU}X^{i}X^{j}+2K_{ij}X^{i}\frac{dX^{i}}{dU}=0.$ (2.2b)
We can immediately solve the $V$-equation (2.2b) if the transverse geodesics $X^{i}(U)$ are known,
$V(U)=-\frac{1}{2}X^{i}\frac{dX^{i}}{dU}+c_{1}U+c_{2}.$ (2.3)
The constants $c_{1}$ and $c_{2}$ are to be determined from the initial conditions. So, the most crucial part of the geodesic problem (thus, of the memory effect) is to solve (2.2a), which, when $K_{ij}$ depends on time, is a Sturm-Liouville problem.
Only a few exact solutions are known, and generally, we confine ourselves to numerical methods (see, e.g, [8]). Analytical or numerical, both types of studies show that geodesics of plane GWs have, generically, a non-vanishing constant velocity after the wave has passed, confirming the VME effect found in [2, 3, 4].
Therefore, it comes as a surprise that the Gaussian profile with just 1 transverse direction
$K^{G}(U)=\frac{k}{\sqrt{\pi}}e^{-U^{2}},$ (2.4)
yields DME if certain critical values are chosen for the amplitude, $k=k_{crit}$ [9, 10]. This confirms the statement of Zeldevich and Polnarev [1], who advocated that the relative velocity will be zero and there will only be a pure displacement at distant times. In that manner, the geodesic should satisfy the Neumann boundary conditions
$\frac{dX^{i}}{dU}|_{U\to\pm\infty}=0,$ (2.5)
and the related equation (2.2a) becomes a Sturm-Liouville (SL) boundary value problem.
In their quest for such magical values $k=k_{crit}$ authors of [9, 10] were motivated to replace Gaussian profile (2.4) with that of Pöschl-Teller (PT)
$K^{PT}(U)=\frac{k}{2\cosh^{2}U},$ (2.6)
which is a good approximation of the Gaussian (2.4).
While for $K^{G}(U)$ (and its derivatives) only numerical solutions are obtained, geodesics of PT profile (2.6) with the condition (2.5) can be found analytically: DME occurs for $k_{crit}=4m(m+1)$ where $m$ is a positive integer. Moreover, their very recent paper [11] reports that the scarf profile
$K^{scarf}(U)=-2g\frac{\sinh U}{\cosh^{2}U}.$ (2.7)
is also a good approximation to flyby [1] described by the derivative of (2.4). Similar to the PT profile (2.6), the Scarf profile is also capable of having a DME if magic parameter values are chosen. In that case, it is even possible to create an exact plane GW with 2-transverse directions, both yielding DME geodesics. This is true also for the derivative of the Gaussian profile.
Let us mention that a PT-type wave profile (2.6) was also employed in [20] to study gravitational wave memories (DME and VME) within Ellis-Bronnikov wormhole geometries. In [21], the same profile was used in Kundt wave spacetimes to indicate a relation between curvature and memory. The authors showed the existence of DME numerically. In [22], similar profiles are adopted to study the so-called ”B-memory” of vacuum plane waves.
Leaving those numerical studies aside, we will focus our attention on analytical/exact results for (2.6) and (2.7). We will treat the related geodesic equations (2.2a) as a non-relativistic Schrödinger equation. The latter is exactly solvable for a few cases. A list of them [5, 6] includes (2.6) and (2.7). Thus, the methods of supersymmetry are directly applicable to the related plane GW geodesics. In that way, we will reveal the origin of DME of plane waves in the context of SUSY QM, which restores the magical values of parameters and necessary conditions naturally. We believe that this similarity of plane wave geodesics with supersymmetry in quantum mechanics is interesting in its own right.
3 Creating plane and gravitational waves via SUSY QM
The transverse part of the geodesic equations in plane waves (2.2a) together with (2.5) constitute the SL boundary value problem [^1] For simplicity, we consider uncoupled equations with vanishing off-diagonal elements.. Therefore, an exactly solvable geodesic equation of this form is expected to have a correspondence in SUSY QM where exact solutions can be derived with its own elegant methods.
Here, we will explicitly reveal the correspondence between geodesics of plane waves in Section 2 and quantum mechanics by deriving the results of [9, 10, 11] with the methods of SUSY. Basic tools of SUSY QM can be recalled from our Appendix A whenever necessary.
Our approach has a couple of advantages:
  1. 1.
    The usage superpotential $W(U)$ (A.1) allows us to unify the recent works about DME of plane GWs in an elegant way.
  2. 2.
    Incorporation of supersymmetric methods explain the magical amplitude values and half waves of DME effect.
  3. 3.
    We can construct further wave profiles with solvable geodesics that can mimic more realistic but not analytically solvable ones.
Let us achieve our first goal.
3.1 Unified wave profiles via SUSY QM
Consider the superpotential
$W(U)=A\text{tanh}(\alpha U)+B\text{sech}(\alpha U).$ (3.1)
This potential is noted in the list of shape invariant potentials in [5, 6] (for earlier studies, see [23, 24]). $A,B,\alpha$ are free parameters for now. Out of this superpotential (3.1), one can derive 2 partner potentials (A.2)
$V_{1}$ $=W^{2}-\frac{dW}{dU}=(B^{2}-A^{2}-A\alpha)\ \text{sech}^{2}(\alpha U)+B(2A+\alpha)\ \text{sech}(\alpha U)\text{tanh}(\alpha U)+A^{2}\qquad\qquad$ (3.2a)
$V_{2}$ $=W^{2}+\frac{dW}{dU}=(B^{2}-A^{2}+A\alpha)\ \text{sech}^{2}(\alpha U)+B(2A-\alpha)\ \text{sech}(\alpha U)\text{tanh}(\alpha U)+A^{2}.\qquad\qquad$ (3.2b)
$V_{1}$ and $V_{2}$ are related with the shape invariance condition (A.5) with $A$ being the shape invariance parameter
$V_{2}(U;A)=V_{1}(U;A-\alpha)+\alpha(2A-\alpha).$ (3.3)
Using the condition (3.3) and shifting $A$, exact bound state spectrum for $V_{1}$ (hence $V_{2}$) can be found
$E_{n}=A^{2}-(A-n\alpha)^{2},\quad n=0,1,2,...$ (3.4)
Above $n=0$ is the ground state with $E_{0}=0$, $n=1$ is the first excited bound state, etc. Please note that their wavefunctions vanish at $U\to\pm\infty$. Below, we will see that their first derivatives are also vanishing at distant times so that they satisfy DME conditions (2.5).
We put the associated Schrödinger equation (A.3) into the form of a geodesic equation (2.2a)
$\frac{d^{2}\psi_{n}}{dU^{2}}-\left[(A-n\alpha)^{2}+(B^{2}-A^{2}-A\alpha)\ \text{sech}^{2}(\alpha U)+B(2A+\alpha)\ \text{sech}(\alpha U)\text{tanh}(\alpha U)\right]\psi_{n}=0.$ (3.5)
Up to a normalization, the ground state wave function can be found (A.4)
$\psi_{0}(U)=e^{-\frac{1}{\alpha}\left(A\ln\big{(}\cosh(\alpha U)\big{)}+B\tan^{-1}\big{(}\sinh(\alpha U)\big{)}\right)}.$ (3.6)
In order to pass to the geodesics (2.2), we make the replacement $\psi(U)\to X(U)$ and set $\alpha=1$ in the above expressions. It will be immediately recognized that (3.5) is composed of 2 parts: PT profile (2.6) in [9, 10] and Scarf profile (2.7) in [11]. Therefore, we have achieved our first goal of a unified framework. It is precisely this unified view that will allow us to understand DME of the related profiles in a systematic way, explaining the quantized values of magical amplitudes.
3.2 DME of plane GWs
Now, it is time for our second goal. Let us begin with $n=1$. Either from (A.7) or (A.8), we get
$X_{1}(U)$ $=$ $L^{\dagger}(U;A)X_{0}(U;A-1),$ (3.7)
$=$ $\operatorname{sech}^{A-1}U\ e^{-B\tan^{-1}(\sinh U)}\left[(2A-1)\tanh U+2B\operatorname{sech}U\right].$
Provided that $A\geq 1$, the velocity of $X_{1}(U)$ is vanishing at infinities
$\frac{dX_{1}}{dU}|_{U\to\pm\infty}=0,$ (3.8)
so that it exhibits DME [^2] It is a half wave in the language of [9, 10, 11]. One can also see that at distant times $U=\pm\infty$, $X_{1}(U)\neq 0$ only if $A=n=1$.
Subsequent SUSY Darboux transformations (A.7) generate other geodesics with higher ”half-wave numbers” $n$.
With this unified approach, we will show below that the PT and Scarf profiles in [9, 10, 11] are actually two different faces of the same theory from the perspective of SUSY QM. Both profiles are rooted in (3.1), hence derivable from (3.2).
One important point is the value of the free parameter $A\geq n$. If we set $A=n$, the formula for geodesics (A.8) simplifies in terms of $W(U)$ as
$X_{n}(U;A=n)=W(U;n)X_{n-1}(U;n-1)-\frac{d}{dU}X_{n-1}(U;n-1).$ (3.9)
The DME is the difference between the asymptotic values of geodesics:
$\Delta X_{n}\equiv X_{n}(U\to\infty;A=n)-X_{n}(U\to-\infty;A=n).$ (3.10)
If we substitute (3.9) into (3.10), we obtain
$\Delta X_{n}$ $=$ $X_{n}(\infty;n)-X_{n}(-\infty;n)$ (3.11)
$=$ $W(\infty;n)X_{n-1}(\infty;n-1)-W(-\infty;n)X_{n-1}(-\infty;n-1),$ (3.12)
as our geodesics satisfy the Neumann boundary condition (2.5). Therefore, SUSY QM yields a nice formula (3.12) for the DME of plane waves in terms of asymptotic values of superpotentials and preceding geodesics333(3.12) might be somewhat reminiscent of Christodoulou’s nonlinear memory [25].. While (3.11) shows that $\Delta X\neq 0$ condition excludes parity even geodesics, second formula (3.12) indicates that DME is related to the parity of superpotential and the preceeding geodesics.
But individual parts of our superpotential (3.1) have different parity properties. Thus, we will investigate each part separately and show that these magical numbers come out of our unified picture in a beautiful manner.
3.2.1 Pöschl-Teller profile
To confine ourselves with the PT type profile (2.6), we need to set
$B(2A+1)$ $=0,$ (3.13a)
$A-n$ $=0$ (3.13b)
in (3.5). While the first condition eliminates the parity-odd scarf term, the second one selects the highest excited state by quantizing the parameter $A$. Selecting $A$ as an integer leads to the so-called reflectionless PT problem. In that case, the geodesic equation turns out to be
$\frac{d^{2}X_{n}}{dU^{2}}+n(n+1)\text{sech}^{2}(U)X_{n}=0,$ (3.14)
cf.(2.2a) As we are dealing with a shape invariant profile, the bounded geodesics can be found via (A.7), where we set $A=m$ first,
$X_{n}(U;m)=L^{\dagger}(U,m)\ L^{\dagger}(U,m-1)...L^{\dagger}(U,m+1-n)X_{0}(U,m-n),$ (3.15)
where $X_{0}(U,m)=\operatorname{sech}^{m}U$ and $m=1,2,3,..$.
Observe that the seed geodesic $X_{0}(U,m)$ is parity even. However, when $B=0$, the superpotential (3.1) becomes parity odd. In addition, each Darboux transformation $L^{\dagger}$ (3.15) changes the parity. Therefore, $\Delta X_{n}$ (3.12) vanishes for even $n$.
For instance, if we take $n=1$, we obtain
$X_{1}(U;m)$ $=$ $L^{\dagger}(U,m)X_{0}(U,m-1),$ (3.16)
$=$ $\left(-\frac{d}{dU}+m\tanh U\right)\operatorname{sech}^{m-1}U=(2m-1)\operatorname{sech}^{m-1}U\tanh U.$
the Neumann boundary condition (2.5) holds for any $m\geq 1$. Further setting $m=n=1$, we get
$X_{1}=\tanh U$ (3.17)
in line with [9, 10]. DME effect can be computed very easily from (3.12) as
$\Delta X_{1}=2.$ (3.18)
So, there is a net displacement, as expected from the parity.
Similarly, we can obtain $n=2$ geodesic using (3.15) as
$X_{2}(U;m)=(2m-3)\Big{(}2(m-1)\operatorname{sech}^{m-2}U+(1-2m)\operatorname{sech}^{m}U\Big{)},$ (3.19)
and for any $m\geq 2$, it satisfies the DME condition (2.5). When $m=2$, the geodesic becomes
$X_{2}=3\tanh^{2}U-1,$ (3.20)
with $\Delta X_{2}=0$ from both (3.11) and (3.12).
For $n=3$, solution of the geodesic equation (3.14) is
$X_{3}(U;m)=(2m-5)(2m-3)\Big{(}2(m-2)\operatorname{sech}^{m-3}U+(1-2m)\operatorname{sech}^{m-1}U\Big{)}\tanh U$ (3.21)
As before, non-vanishing displacement occurs only when $m=n=3$, such that
$X_{3}(U;m=3)=6\tanh U-15\tanh U\operatorname{sech}^{2}U,\quad\Delta X_{3}(m=3)=12.$ (3.22)
One can go on and verify that $\Delta X_{n}(m)\neq 0$ only for $m=n$ and $n$ is an odd integer. We conclude that the fine-tuned parameters yielding DME are nothing but the highest excited states of the analogous quantum mechanical problem satisfying the Neumann boundary conditions.
Longitudinal motion and memory
In order to compute $V$-component of geodesics for the first two $n$, we substitute our solutions for $X(U)$ i.e., (3.17) and (3.20) into (2.3) respectively. We observe that part of the longitudinal motion organizes itself as another particular solution of the PT problem. For instance,
$V_{1}(U)$ $=$ $-\frac{1}{2}\tanh U\operatorname{sech}^{2}U+c_{1}U+c_{2}=-\frac{1}{10}X_{1}(U;m=3)+c_{1}U+c_{2},$ (3.23)
$V_{2}(U)$ $=$ $-3\tanh U\operatorname{sech}^{2}(2-3\operatorname{sech}^{2}U)+c_{3}U+c_{4}=-\frac{1}{35}X_{3}(U;m=5)+c_{3}U+c_{4},\qquad$ (3.24)
where $c_{1},..,c_{4}$ are constants. $X_{1}(U;m=3)$ and $X_{3}(U;m=5)$ can be found from (3.16) and (3.21), respectively. As $n\neq m$, those motions exhibit VME provided that initial constants $c_{1},c_{3}$ do not vanish.
3.2.2 Scarf profile
In order to reproduce the geodesic equation in [11] for scarf profile (2.7), we need to satisfy another set of equations
$A-n$ $=0,$ (3.25a)
$B^{2}-A^{2}-A$ $=0,$ (3.25b)
in (3.5). While the first equation quantizes the parameter $A$, the second one solves for $B$ as
$B=\pm\sqrt{n(n+1)},\quad A=n.$ (3.26)
where $n=0,1,2,\cdots$. Pay attention to the fact that $B$ has double roots. This multiplicity will allow us to extend our arguments to exact plane gravitational waves with 2-transverse directions [^4] That extension was first done in [11]. Here, we elaborate on their results using SUSY methods. in Section 4.
In $1$-transverse dimensional case, (3.5) reduces to a geodesic equation (2.2)
$\frac{d^{2}X_{n}}{dU^{2}}\mp(2n+1)\sqrt{n(n+1)}\ \text{sech}U\ \text{tanh}U\ X_{n}=0,$ (3.27)
augmented with (2.5). Thus, supersymmetry leads to a very quick derivation of the scarf geodesic equation that mimics the derived Gaussian flyby profile. Note that, in [11] above equation was obtained through the hypergeometric equation and solved with the Nikiforov-Uvarov method. SUSY provides a more economical way for the solutions.
From the viewpoint of quantum mechanics, $X_{n}(U)$ is not the zero-energy ground state of the scarf potential (2.7). Rather, it corresponds to a particular parametrization of its parameters $A,B$ with the quantum number $n$ (3.26). Let us derive DME within this profile.
The seed geodesic can be found via (A.4) as
$X_{0}=\operatorname{sech}^{m}U\ e^{-B\tan^{-1}\sinh U},$ (3.28)
where $A=m$ and $B=\pm\sqrt{n(n+1)}$. As before, $m$ is a positive integer. When $n=1$ geodesic can be obtained from (A.7)
$X_{1}(U;m)=\left[(2m-1)\tanh U+2B\operatorname{sech}U\right]\operatorname{sech}^{m-1}U\ e^{-B\tan^{-1}(\sinh U)}.$ (3.29)
It is only when $m=n=1$ (3.29) yields a non-vanishing asymptotic value. In this case, it reduces to
$X_{1}(U;m=1)=\left[\tanh U+2B\operatorname{sech}U\right]e^{-B\tan^{-1}(\sinh U)}.$ (3.30)
Let us emphasize that (3.30) satisfies (3.27) with $n=1$. The related DME can be computed either directly from the geodesic solution or from (3.12)
$\Delta X_{1}=2\cosh\left(\frac{B\pi}{2}\right)=2\cosh\left(\frac{\pi}{\sqrt{2}}\right),\quad B=\pm\sqrt{2}$ (3.31)
which is independent of the sign of $B$.
Applying a second Darboux transformation, we generate $X_{2}(U;m)$. In order to satisfy (3.27), we immediately set $m=2$ and get
$X_{2}(U;m=2)=\left[2+21\operatorname{sech}^{2}U+8B\tanh U\operatorname{sech}U\right]\ e^{-B\tan^{-1}(\sinh U)},$ (3.32)
with $B^{2}=6$. The DME of this second geodesic is sensitive to the sign of $B$
$\Delta X_{2}=-4\sinh\left(\frac{B\pi}{2}\right),$ (3.33)
as $B=\pm\sqrt{6}$.
By virtue of supersymmetric transformations (3.9), one may obtain geodesics with non-zero DME in a straightforward manner.
Longitudinal motion and memory
In order to illustrate the longitudinal memory, we may construct $V_{1}(U)$ (2.3)
$V_{1}(U)$ $=$ $-\frac{1}{2}X_{1}(U;m=1)\frac{dX_{1}(U;m=1)}{dU}+c_{1}U+c_{2},$ (3.34)
$=$ $-3\operatorname{sech}U\Big{(}5\tanh U\operatorname{sech}U+B(1+\operatorname{sech}^{2}U)\Big{)}e^{-2B\tan^{-1}(\sinh U)}+c_{1}U+c_{2}.\qquad$ (3.35)
As
$\frac{dV_{1}}{dU}|_{\pm\infty}=c_{1},$ (3.36)
where the coordinate $V$ exhibits a VME. This result is generic for any $V_{n}(U)$.
4 4-dimensional plane waves
Apart from those 3-dimensional cases presented above, we may construct plane waves with 2-transverse directions [^5] One may try combinations of Scarf and PT profiles too.. For this, we choose a diagonal profile $K$ in (2.1), set $X^{i}=\{X,Y\}$ and begin with Scarf-type profile first.
4.1 Exact plane gravitational wave with Scarf profile
The double root of $B$ (3.26) allows Zhang et al. [11] to construct an exact plane gravitational wave (vacuum solution of Einstein’s equations) mimicking the derivative Gaussian profile (2.4). The associated profile is diagonal
$K_{scarf}=\begin{pmatrix}(2n+1)|B|\operatorname{sech}U\tanh U&0\\ 0&-(2n+1)|B|\operatorname{sech}U\tanh U\,\end{pmatrix}$ (4.1)
and traceless. $|B|=\sqrt{n(n+1)}$ and $n=1,2,3,...$.
First, 2 sets of geodesics with their DMEs can be simply extracted from Section 3.2.2 as
$X_{1}(U)$ $=\left[\tanh U+2\sqrt{2}\operatorname{sech}U\right]e^{-\sqrt{2}\tan^{-1}(\sinh U)},\quad\quad\Delta X_{1}=2\cosh\left(\frac{\pi}{\sqrt{2}}\right),$ (4.2a)
$Y_{1}(U)$ $=\left[\tanh U-2\sqrt{2}\operatorname{sech}U\right]e^{\sqrt{2}\tan^{-1}(\sinh U)},\quad\quad\Delta Y_{1}=2\cosh\left(\frac{\pi}{\sqrt{2}}\right),$ (4.2b)
$X_{2}$ $=\left[2+21\operatorname{sech}^{2}U+8\sqrt{6}\tanh U\operatorname{sech}U\right]e^{-\sqrt{6}\tan^{-1}(\sinh U)},\quad\Delta X_{2}=-4\sinh\Big{(}\frac{\sqrt{6}\pi}{2}\Big{)},$ (4.3a)
$Y_{2}$ $=\left[2+21\operatorname{sech}^{2}U-8\sqrt{6}\tanh U\operatorname{sech}U\right]e^{\sqrt{6}\tan^{-1}(\sinh U)},\quad\Delta Y_{2}=4\sinh\Big{(}\frac{\sqrt{6}\pi}{2}\Big{)}.$ (4.3b)
The above results are in line with the ones in [11].
It is also possible to find lightcone geodesic $V_{n}(U)$ by combining solutions of transverse geodesics. As explained in Section 3.2.2, the longitudinal memory will be a VME.
We may construct further plane wave solutions where $m>n$, see (3.29). In that case, the trace of the profile is non-zero, i.e., $\mathrm{Tr}{K}=2(m-n)^{2}$. One can get a non-vanishing DME when, for instance, $n=2$ and $m=3$.
4.2 Pöschl-Teller plane waves
In [10], it was argued that when an exact plane GW is built out of a PT profile, DME can be obtained only in one direction. The other direction should be put to zero by initial conditions. SUSY provides a reasoning for this: Unlike the scarf case, the parameter $A$ of the PT profile has a single root (3.13b).
To understand this better, we may construct a 4-dimensional plane wave out of supersymmetric partners (3.2). The result will be a null-fluid spacetime, $\mathrm{Tr}(K)\neq 0$, with the following diagonal profile
$K_{PT}=\begin{pmatrix}(m-n)^{2}-m(m+1)\operatorname{sech}^{2}U&0\\ 0&(m-n)^{2}-m(m-1)\operatorname{sech}^{2}U\,\end{pmatrix}$ (4.4)
where parameters $m$ and $n$ are integers. From Section (3.2.1), we recall that DME occurs only when $m=n$. That simplifies (4.4) and puts it into its final form. Then, resulting geodesic equations (2.2a) are
$\left(\frac{d^{2}}{dU^{2}}+m(m+1)\operatorname{sech}^{2}U\right)X$ $=0,$ (4.5a)
$\left(\frac{d^{2}}{dU^{2}}+m(m-1)\operatorname{sech}^{2}U\right)Y$ $=0$ (4.5b)
Due to shape invariance, the solutions of $X$ and $Y$ are interrelated: they match with each other except the ground state of $X$. If one sets $m=1$, the solution for the first one is $X=\tanh U$ (3.17). The second equation, on the other hand, becomes that of a free particle and confirms the claim made in [10]. Recall that choosing $m$ as an integer leads to a ”reflectionless PT potential” whose reflection coefficient is vanishing just like a free particle.
Other choices of $m>1$ yield DME again in one direction. For instance, when $m=2$ $X=3\tanh^{2}U-1$ and $Y=\tanh U$. Therefore, the usage of superpartners in the profile yields a tower of geodesics exhibiting DME in alternating directions.
5 Modeling singular profiles with SUSY
It is time for our third goal. We show that SUSY provides further insight and predicts DME for other GWs.
In [26, 27], memory effects of the profile $K=\frac{1}{U^{4}}$ was discussed. Later, authors in [7] obtained a 1-directional DME within the same singular plane GW.
There are also exact plane gravitational waves with a $\frac{1}{U^{2}}$ singularity near the origin. The simplest example is the one with the profile $K=\frac{k}{U^{2}}\rm{diag}(1,-1)$ whose symmetries worked in [28]. There is also the Lukash gravitational wave [29, 30], which is endowed with an extra symmetry but suffers from the same singularity [^6] Same singularity appears in nuclear physics, see [31] and references therein..
Now, we show that SUSY QM allows us to seek DME within such singular waves. The approximated profile can be constructed out of the superpotential
$W(U)=A\coth U,$ (5.1)
which is singular at the origin. Partner potentials, in this case, are
$V_{1}$ $=A^{2}+A(A+1)\operatorname{csch}^{2}U;$ (5.2a)
$V_{2}$ $=A^{2}+A(A-1)\operatorname{csch}^{2}U.$ (5.2b)
This singular potential is shape invariant with $A\to A-1$; thus integrable. Note that $W(U)$ (5.1) is parity odd; e.g., $W(\infty)=-W(-\infty)=1$. In order to obtain a non-vanishing DME, we select $A=n$, where $n$ indexes the bound states. Then, the associated geodesic equation becomes
$\left(\frac{d^{2}}{dU^{2}}-n(n+1)\operatorname{csch}^{2}U\right)X=0,$ (5.3)
and its solutions can be immediately found with the technique summarized in Section 3.2. For instance,
$X_{1}$ $=\coth U;\qquad\qquad\ \ \Delta X_{1}=2;$ (5.4a)
$X_{2}$ $=3\coth^{2}U-1;\qquad\Delta X_{2}=0.$ (5.4b)
These solutions blow up at $U=0$, as expected. Using the parity argument given in (3.11), we deduce that only geodesics indexed with odd $n$ have non-vanishing DME. For instance, we obtain $\Delta X_{3}=12$ without even solving for $X_{3}(U)$.
If we were to build a plane wave including the partner potential $V_{2}$, it would be pretty much similar to the PT plane wave described in Section (4.2). On the other hand, if we build an exact plane GW with $K$ being
$K_{sing}=\begin{pmatrix}n(n+1)\operatorname{csch}^{2}U&0\\ 0&-n(n+1)\operatorname{csch}^{2}U\,\end{pmatrix}$ (5.5)
where we obtain DME only in the $X$-sector.
Based on our model, we conclude that DME can also be found in GWs endowed with $1/U^{2}$ singularity. Of course, memory effects in singular profiles should be considered with more care. Our arguments here can be taken as a necessary first step towards a complete analytic solution.
6 Discussion
In this work, based on the idea that DME conditions in plane GWs match the ones in quantum mechanics, we rederive the recent results of [9, 10, 11] within the supersymmetric framework. However, our work extends a mere rederivation in several ways. First of all, SUSY methods unify the Scarf and Pöschl-Teller GW profiles in an elegant way, showing that they are two different reductions of the same theory. The reduction process amounts to satisfying a couple of equations, which automatically restores the amplitudes of the wave profiles: no fine-tuning is necessary.
Geodesics with DME, i.e., $\Delta X\neq 0$, correspond to certain bound states of analogous quantum problems. Even in that case, the explanation depends on the parity and the quantization of the parameter in a certain way. Thanks to the shape invariance property of SUSY QM, solutions for the geodesic motion can be found readily using Darboaux transformations. For the PT profile, DME occurs with odd $n$. For the Scarf profile, DME occurs for all $n$. See Table 1 for some of our results.
Table 1: SUSY generated DME
W(U) Profile Quantized parameters DME ($\Delta X_{1}$) Parity
$A\tanh U+B\operatorname{sech}U$ Pöschl-Teller $A=n,\ B=0$ $2$ even
$A\tanh U+B\operatorname{sech}U$ Scarf $A=n,\ B=\pm\sqrt{n(n+1)}$ $2\cosh(\frac{\pi}{\sqrt{2}})$ odd
$A\coth U-B\operatorname{csch}U$ Singular $A=n,\ B=0$ $2$ even
In addition, we propose a hypothetical 2-d transverse PT plane wave, which explains its half DME in itself. $4$-dimensional Scarf exact GW fits to our framework because of the $\pm$ ambiguity in its quantized parameter $B$.
Lastly, we make a first attempt to study DME in GWs with $\frac{1}{U^{2}}$ singularity. Our SUSY motivated model predicts the existence of the DME, see Section 5. Because of singularity, this investigation requires much care, and we reserve its complete treatment for a follow-up study where we also plan to work out DME in cosmological Lukash GW.
Another research problem is to look for DME in the Kepler GW [32], which corresponds to the Coulomb potential in quantum mechanics. Thanks to the integrability, we know the solutions of the Sturm-Liouville problem which may allow us to go to BJR coordinates and write down the Carroll symmetry generators at once.
There are further questions to be explored. The list of SUSY QM provides us with other exactly solvable potentials. Among them, there are trigonometric Scarf and trigonometric Pöschl-Teller potentials which can be prototypes for periodic GWs. We would like to study DME of periodic GWs using them. Another interesting direction is the double copy of the Scarf and the PT profiles that we worked out. We wonder what kind of singular electromagnetic configurations they yield and what their helicities will be.
As we know, the Pöschl-Teller equation frequently occurs in soliton physics. We would like to understand whether there is a relation between the DME of PT profile and the topology of solitons. Likewise, it would be interesting to understand the role of the Schwarzian derivative in [12] in the Scarf case. Finally, there is a recent memory type called the frequency memory effect [21] in plane GWs. We can also study such effects in our SUSY framework in the future.
Acknowledgments
We are grateful to Krzysztof Andrzejewski, Peter Horvathy and Zurab Silagadze for their comments on our work.
M. E. is supported by İstanbul Bilgi University research fund (BAP) with grant no: 2024.01.009 and by The Scientific and Technological Research Council of Turkey (TÜBİTAK) under grant number 125F021. M. E. is also funded partially by the TÜBİTAK BİDEB 2232-A program under project number 121C067. M. M. is supported by the Scientific and Technological Research Council of Turkey (TÜBİTAK) under grant numbers 122F451 and 123N952. E. C. is partially funded by the Scientific and Technological Research Council of Turkey (TÜBİTAK) under grant number 125F021.
Appendix A SUSY QM: A brief review
Below, we provide a brief summary of supersymmetric methods in QM to be self-consistent. We suggest [5, 6] for more details.
We begin with the main elements of SUSY QM, namely its superpotential $W(U)$ and ladder operators $L,L^{\dagger}$ such that
$L=\frac{d}{dU}+W(U),\quad L^{\dagger}=-\frac{d}{dU}+W(U).$ (A.1)
They factorize Hamiltonians
$H_{1}$ $=L^{\dagger}L=-\frac{d^{2}}{dU^{2}}+V_{1},\quad\quad V_{1}=W^{2}-\frac{dW}{dU},$ (A.2a)
$H_{2}$ $=LL^{\dagger}=-\frac{d^{2}}{dU^{2}}+V_{2},\quad\quad V_{2}=W^{2}+\frac{dW}{dU},$ (A.2b)
where $V_{1,2}$ are partner potentials sharing almost the same spectrum, except the zero-energy ground state of $V_{1}$. SUSY also relates the scattering states of that partner potentials. If the potential is shape invariant, the exact spectrum can be found easily. It is this crucial property of shape invariance of SUSY QM that will allow us to derive DME of plane GWs.
The time-independent Schrödinger equation for the first potential $V_{1}$ is
$\left(-\frac{d^{2}}{dU^{2}}+V_{1}\right)\psi_{n}=E_{n}\psi_{n},$ (A.3)
where $n=0,1,2,...$ indexes the bound state. Up to a normalization factor, zero energy bound state $\psi_{0}(U)$ can be found via
[…]
immediately.
If partner potentials $V_{1,2}$ are shape invariant, i.e. connected via
$V_{2}(U;A_{1})=V_{1}(U;A_{2})+f(A_{1}),$ (A.5)
where $A_{1,2}$ are parameters [^7] Recall that plane wave profiles (2.4), (2.6) and (2.7) also have parameters like $k,g,..$., energy eigenvalues and eigenstates of (A.2) can be computed easily. Similar to $f(A_{1})$, $A_{2}$ is a function of the first parameter as well, i.e., $A_{2}=A_{2}(A_{1})$. Due to (A.5), discrete eigenstates are found step by step
$\psi_{n}(U;A_{1})=L^{\dagger}(U;A_{1})\psi_{n-1}(U;A_{2}).$ (A.6)
For instance, if $A_{1}$ and $A_{2}$ are related by a simple relation $A_{1}-A_{2}=1$, the excited state wave functions can be computed with subsequent action of $L^{\dagger}$ (A.1)
$\psi_{n}(U;A)=L^{\dagger}(U;A)L^{\dagger}(U;A-1)...L^{\dagger}(U;A-n+1)\psi_{0}(U;A-n),$ (A.7)
where we set $A_{1}=A$. In that case, (A.6) simplifies as
$\psi_{n}(U;A)=\left(-\frac{d}{dU}+W(U;A)\right)\psi_{n-1}(U;A-1).$ (A.8)
For now, the parameter $A$ can take any value. It has a physical meaning when applied to geodesics at Sections 3, 4, and 5.